Suppose b,c 2 R, and T: R3 ! R2 deÖned as

T (x; y; z) = (2x x 4y + 3z + b; 6x + cxy):

Show that T is linear if and only if b = c = 0


1
Expert's answer
2021-06-17T16:00:47-0400
"T (x;y;z) = (2x-4y+3z+b,6x+cxy)"

Let us consider "T(x_1+x_2,y_1+y_2,z_1+z_2)." It should be equal to 

"T(x_1,y_1,z_1) + T(x_2,y_2,z_2)"

 .


"T(x_1+x_2,y_1+y_2,z_1+z_2)"




"= \big(2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+b,"

"6(x_1+x_2)+c(x_1+x_2)(y_1+y_2)\big)."



"T(x_1,y_1,z_1) + T(x_2,y_2,z_2)"




"= (2x_1-4y_1+3z_1+b,6x_1+cx_1y_1)"

"+ (2x_2-4y_2+3z_2+b,6x_2+cx_2y_2 )"

"= \big(2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+2b,"


"6(x_1+x_2)+c(x_1y_1+x_2y_2)\big)."

Therefore, 

"2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+b"

 

"= 2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+2b"

for every "x_1,x_2,y_1,y_2,z_1,z_2\in\R." It is true if and only if

 

"2b=b \; \Rightarrow b = 0."


Also it should be true that


"6(x_1+x_2)+c(x_1+x_2)(y_1+y_2)"

 

"=6(x_1+x_2)+c(x_1y_1+x_2y_2)"




"c(x_1+x_2)(y_1+y_2)=c(x_1y_1+x_2y_2)"


"c(x_1y_2+x_2y_1)=c"

for every "x_1,x_2,y_1,y_2,z_1,z_2\in\R." 

Let, for example, "x_1=y_1=x_2=y_2 = 1,"  then


"2c=c=>c = 0."

 So "c" can be only equal to 0.


So we conclude that if T is linear, "b=0" and "c=0."


Now we should prove that for "b=c=0" "T" is linear.



"T(x,y,z) = (2x - 4y + 3z, 6x)"



For every "x_1,x_2,y_1,y_2,z_1,z_2\in\R \; \;" 


"T(x_1+x_2,y_1+y_2,z_1+z_2)"




"= (2(x_1+x_2)-4(y_1+y_2)+3(z_1+z_2),6(x_1+x_2))"

"= (2x_1-4y_1+3z_1,6x_1) + (2x_2-4y_2+3z_2,6x_2)"

"= T(x_1,y_1,z_1) + T(x_2,y_2,z_2),"

so "T" is linear.